OpenStudy (anonymous):

Find the 9th term of the geometric sequence with a_2 0.3 and a_6= 0.00003

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

geometric sequence means you get the next term by multiplying the previous term by a specific number "x" a_3= a_2 *x a_4= a_3 *x =a_2 *x^2 a_5= a_2 * x^3 a_6= a_2 * x^4 since you know a_6 and a_2 and can solve for x